60x-120-22.5+10x-60+105=24x-54+27.3+20x-120
Answer by fractalier(6550)

Let's collect like terms on both sides...
60x-120-22.5+10x-60+105=24x-54+27.3+20x-120
becomes
70x - 97.5 = 44x - 146.7
Now subtract 44x and add 97.5 to both sides and get
26x = -49.2
and divide by 26 and get
x = -49.2/26 = -1.89 or so